# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v3

@@ Abstract
Index Terms:
real-time, 3D human reconstruction, single image, monocular 3D reconstruction.
@@ I Introduction
Reconstructing a 3D human from a single image has emerged as a popular task in computer vision and graphics, which can be widely used in various downstream applications, such as mixed reality and virtual try-on. However, real-time, high-fidelity monocular 3D human reconstruction remains challenging. The core of this challenge lies in the 3D representation, as it significantly influences the design and performance of reconstruction and generation approaches. Despite some promising results, existing methods [1, 2, 3, 4] typically suffer from high computational consumption and lack of robustness. Consequently, a good 3D geometry representation is essential for 3DTV, Holographic Telepresence systems, and other real-time applications, which must fulfill the requirements of accuracy, efficiency, and compatibility [5].
@@
Classic representations, such as voxel grids [6] or meshes [7, 8, 9], have been explored for monocular human reconstruction. However, voxel grids require a space complexity of $O(n^{3})$ (where $n$ is the resolution of grids in each dimension), and meshes struggle with topology changes or large deformations. Although some recent methods [4, 10] propose optimization-based pipelines combining normal maps, voxel grids, and meshes, they all suffer from efficiency issues.
Implicit neural representations have emerged and been widely used in monocular human reconstruction [1, 11, 12, 13, 3]. These methods treat 3D space as a continuous field, such as an occupancy or signed distance field, represented as $F(x,y,z):\mathbb{R}^{3}\to\mathbb{R}$. Instead of voxel grids, a neural network models the field, allowing results at any resolution. However, these methods infer values across numerous spatial grid points, which slows down inference and makes high-frame-rate or real-time reconstruction a significant challenge. Based on PIFu [1], Monoport [14] proposes an efficient sampling scheme to speed up inference, achieving 15 FPS only for mesh-free rendering. However, when generating a complete mesh, their method operates at less than 10 FPS, and the results are often worse than PIFu. These limitations hinder Monoport’s applicability with downstream applications.
@@
Parametric Model [15, 16, 17]
Voxel Grid[6]
Mesh[8, 9, 7]
Implicit Function[1], [11, 12, 13, 3, 18]
@@
  • Compared with state-of-the-art methods based on other representations, our approach can produce high-quality results in real-time. Our system is the first 30+FPS pipeline, achieving a twofold improvement in speed over Monoport [14], coupled with superior quality.
An early version of our method, including only FOF, was published as a conference paper [19]. Based on our FOF representation, several works for generation[20] and reconstruction[21, 22, 23] have been proposed. In this paper, we substantially extend the original version by developing FOF-X. Firstly, we design a new reconstruction framework that greatly mitigates the performance degradation caused by texture and lighting effects. Secondly, we propose a robust mesh-to-FOF conversion algorithm with an automaton-based discontinuity matcher, enabling real-time execution and significantly improving the system’s robustness when facing challenging poses. Thirdly, we propose a FOF-to-mesh algorithm with a Laplacian coordinate constraint for greater robustness and fidelity. Such a strategy effectively addresses artifacts caused by view-direction bias without any loss of geometric details. Fourth, we employ cosine series as the subspace approximation basis, retaining the Fourier-like approximation capacity while eliminating Gibbs phenomena through even-periodic boundary extension. Additionally, we evaluate FOF-X through comprehensive experiments including: testing on more datasets, comparisons with various state-of-the-art methods, and more detailed ablation studies, all confirming the method’s effectiveness, efficiency and robustness.
@@ II Related Work
Regarding geometric representations, we categorize mainstream monocular 3D human reconstruction approaches into three types: surface-based reconstruction, volume-based reconstruction, and “sandwich-like” reconstruction. Note that “sandwich-like” reconstruction can be considered a blend of surface-based and volume-based methods. Additionally, generative models can also be used directly or indirectly for single-image 3D human reconstruction. We review these methods in this section.
@@ II-A Surface-based Reconstruction
Surface-based approaches focus on inferring the interface between geometry and empty space, and can be broadly classified into three categories: parametric models, UV-map-based methods, and graph-based methods.
Parametric models, such as SMPL [15], SMPL-X [16], and STAR [17], are popular representations for naked 3D humans. These models are derived from large datasets of naked human shapes and use statistical methods to create analytical models, which can generate a human body mesh with just a few dozen parameters. Building on these models, many works [24, 25] estimate 3D human shapes from an RGB image by predicting the parameters of the parametric model. However, these methods are limited in that they cannot reconstruct clothing or hair.
UV-map-based methods attempt to estimate clothed humans by adjusting the vertices of the parametric models. Alldieck et al[8] warp the input image to align with the UV map and calculate the displacements on the T-pose SMPL mesh. However, their results are always in T-pose and do not align well with the input images. Zhu et al[26, 7] employ a four-stage process constrained by joints, silhouettes, and shading, but their results are sometimes inconsistent with the images.
Graph-based representations are naturally compatible with triangle meshes. Li et al[9] use graph neural networks to reconstruct topology-consistent geometries of clothed humans, but the recovered mesh tends to be smooth and lacks fine details. Representing 2D manifold surfaces of different 3D objects with a fixed structure is very challenging. Consequently, most methods heavily rely on the topology defined by SMPL and cannot produce detailed human geometries with complex topology. Habermann et al[27] propose a deep learning approach for monocular dense human performance capture, but this method requires scanning the actor with a 3D scanner to create a template mesh. In a word, graph-based methods struggle to align image features accurately with the results, leading to unstable and over-smoothed outputs.
@@ II-B Volume-based Reconstruction
Volume-based methods estimate attributes such as occupancy and signed distance for each point in the 3D space, allowing them to represent 3D shapes with arbitrary topology. Voxel grids and implicit neural networks are the two main representations used in volume-based methods.
Zheng et al[6] regress voxel grids of 3D humans using convolutional neural networks, but this approach requires intensive memory and results in models with limited resolution. Unlike voxel grids, implicit neural representations can depict detailed 3D shapes without resolution limitations. Saito et al[1] use pixel-aligned functions to reconstruct a 3D human from a single RGB image and further improve 3D geometric details using normal maps [11]. Zheng et al[12] use a voxelized SMPL mesh as a prior to enhance robustness. Xiu et al[3] propose a method to correct the SMPL model estimated by other methods [24, 25, 28] based on the input image, and then regress the occupancy field from the normal map and the signed distance field of the corrected SMPL. While their results are robust to human poses, the geometric details often do not match the input images and can be noisy.
Further advancements like GTA [29], SiFU [2], and D-IF [30] explore implicit neural representations with a globally-correlated 3D-decoupling transformer and uncertainty-aware implicit field inference, respectively. However, they still struggle with computational efficiency. Although Li et al[14] propose an efficient sampling scheme to speed up inference, they achieve 15 FPS only for mesh-free rendering, and when a complete mesh needs to be generated, their method operates at less than 10 FPS.
Recent advancements have seen the utilization of Neural Radiance Fields (NeRF) [31] in various reconstruction approaches, and subsequent methods [32, 33, 34] have been further adapted for human reconstruction. NeRF applies volume rendering to implicit volume-based representations, enabling direct supervision of 3D scenarios with 2D images. Consequently, SHERF [33] achieves the first generalizable Human NeRF from a single human image, while ELICIT [34] employs a pre-trained CLIP model [35] for contextual understanding. However, NeRF-based approaches primarily focus on view synthesis and often lack detailed geometry generation from monocular images.
@@
In recent works [36, 37, 4], 3D objects are modeled using front and back depth maps, resembling a sandwich structure that encloses the inner space. However, depth maps only represent part of the geometric surface and cannot fully reconstruct the object. Therefore, these methods require volume-based post-processing, such as IF-Nets [38] and Poisson surface reconstruction[39], to complete the result further. Any-Shot GIN [37] extends the sandwich-like scheme to novel classes of objects by predicting the depth maps of two sides from a single image and performing shape completion with IF-Nets. ECON [4] follows a similar approach and extends it to 3D humans. It estimates the normal maps of two sides and converts them into depth maps using an optimization method. The parametric human model prior is also employed in the shape completion stage to “inpaint” the missing geometry. However, this optimization process is time-consuming, and its instability may lead to incorrect geometric estimates.
@@ II-D Leveraging Generative Models
Reconstructing a 3D human body from a single image is inherently challenging due to incomplete observations, making it difficult to regress geometric information accurately. In this case, generative models can offer plausible completion for invisible regions. SiTH [18] generates the back-view image with an image-conditioned diffusion module and then predicts dual-sided normal maps for subsequent reconstruction. TeCH [10] recovers human meshes with a SMPL-X initialized hybrid Deep Marching Tetrahedra (DMTet) representation and Score Distillation Sampling (SDS) Loss [40], achieving visually impressive results but with noisy surfaces. Human-LRM [41] first uses an LRM [42] model to reconstruct the human body as a NeRF from a single image. It then applies a diffusion model to refine the multi-view images rendered from the NeRF, followed by another LRM reconstruction based on the refined images to produce the final result, but does not provide a substantial improvement in reconstructed geometry quality. PSHuman [43] first uses a multi-view image diffusion model to generate six views of global full-body images and local face images. Then, using the generated normal and color maps, SMPL-X is deformed and remeshed by differentiable rasterization. However, the heavy pipeline makes it inefficient for real-time use. These generative model-based methods are typically time-consuming, and the reconstructed result does not accurately match the input image.
To solve the problems with existing representations, we present Fourier Occupancy Field (FOF) in this paper. FOF is well-aligned with the input image and offers superior efficiency compared to existing representations. Leveraging the proposed efficient and flexible 3D geometry representation of FOF, we design an extensive reconstruction framework FOF-X and parallelized inter-conversion algorithms between FOF and meshes. Based on FOF-X, we contribute a robust high-quality human reconstruction system operating at 30 FPS with a single RGB camera.
@@ III-A1 Vanilla FOF
An overview of our FOF representation is shown in Fig. 4. Without loss of generality, 3D objects to be reconstructed are normalized into a $1,1]^{3}$ cube. Denoted by $(x,y,z)\in[-1,1]^{3}$ the 3D coordinates, and by $S$ the surface of a 3D object, the 3D object can be described by a 3D occupancy field $F:[-1,1]^{3}\mapsto\{0,0.5,1\}$, where the occupancy function $F(x,y,z)$ is defined as
@@
Such a 3D representation is highly redundant, because only a small subset, i.e., the iso-surface of $F(x,y,z)$ with value of 0.5, is sufficient to represent the surface of the object. Direct inference of the 3D occupancy field $F$ from a single image not only confronts with the curse of dimensionality, but also requires more computation and memory in handling high dimensional feature maps. Note that the occupancy field defined in the 3D cube can be regarded as the collection of a large number of 1D signals defined on the lines along the view direction. Without loss of generality, we assume that the view direction is the same as the $z$-axis of the 3D cube. The occupancy function, as a 1D signal $f(z):[-1,1]\mapsto\left\{0,0.5,1\right\}$, along the line passing through a particular point $(x^{*},y^{*})\in[-1,1]^{2}$ on the xyxy-plane can be written as $f(z)=F(x^{*},y^{*},z)$. We explore Fourier representation for such a family of 1D occupancy signals and propose a more compact 2D vector field orthogonal to the view direction, namely Fourier Occupancy Field (FOF), for more efficient representation of the 3D occupancy field, which is detailed in the following subsections.
Fourier Series on a Single Occupancy Line: 1D occupancy signals $\left\{f(z)\right\}$ are essentially on-off signals switching at the boundaries of human bodies, which lie in a low-dimensional manifold in the ambient signal space. For accurate and fast 3D reconstruction, the compact representation should have the following merits:
  • Sampling Scalability: The representation should be able to adapt to different sampling rates in the inference stage without introducing systematic mismatches except for the inherent approximation error due to sampling. Without retraining, systems with such a representation would support diverse applications with different resolution, quality, and speed requirements.
  • Stable Reconstruction: The representation should provide stable reconstruction quality for diverse humans with various parameters such as garments, heights, weight, and poses.
  • Low-complexity Reconstruction: The representation should allow low-complexity reconstruction to support emerging real-time applications such as Holographic Telepresence.
However, most popular representations do not enjoy all the above merits. For instance, discrete Fourier transform and discrete wavelet transform have stable and low-complexity reconstruction, but only support the sampling rate used in preparation of the training set due to their discrete nature. Learned dictionaries are able to provide sparse representations, but require computation-demanding iterative reconstruction, leaving alone the sampling scalability. Without a doubt, sampling scalability is the most difficult property to meet. Pushing the concept of sampling scalability to the limit, one representation would be scalable to arbitrary sampling rates in theory if it allows continuous reconstruction along the $z$-axis. Note that in practice the representation should stay discrete and compact for efficient inference. Such a bridge of discrete-representation and continuous-reconstruction motivates us to use Fourier series for the representation of 1D occupancy signals $\left\{f(z)\right\}$. By nature, the Fourier series would have an infinite number of discrete coefficients, which is impractical. Note that most energy of such a 1D occupancy signal concentrates on only a few low-frequency terms, which yields a compact representation by subspace approximation.
Formally, let $f_{p}(z)$ be periodic extension of $f(z)$. Note that $f_{p}(z)$ satisfies the Dirichlet conditions:
  • $f_{p}(z)$ is absolutely integrable over one period;
  • $f_{p}(z)$ has a finite number of discontinuities within one period;
  • $f_{p}(z)$ has a finite number of extreme points within one period.
Thus, $f_{p}(z)$ can be expanded as a convergent Fourier series:
@@
where $\{a_{n}\},\{b_{n}\}\in\mathbb{R}$ are coefficients of basis functions $\left\{\cos(nx)\right\}$ and $\left\{\sin(nx)\right\}$, respectively. Note that by defining the occupancy function as Eq. (1), Eq. (2) holds for discontinuity at $z^{*}$ on the surface $S$ because $f(z^{*})=(f(z^{*}-)+f(z^{*}+))/2=0.5$, where $f(z^{*}-)$ and $f(z^{*}+)$ are the left hand limit and right hand limit of $f(z)$ at $z^{*}$. For compact representation, we approximate 1D occupancy signals by a subspace spanned by the first $2N+1$ basis functions:
@@
Fourier Occupancy Field: Such a Fourier subspace approximation is applied to all 1D occupancy signals over the xyxy-plane:
@@
where $\bm{C}(x,y)$ is the ($2N+1$)-dimensional Fourier coefficient vector for the 1D occupancy signal at $(x,y)$. In this way, we obtain the Fourier Occupancy Field $\bm{C}:[-1,1]^{2}\mapsto\mathbb{R}^{2N+1}$ for 3D occupancy field $F$.
Note that 3D coordinates $(x,y,z)$ in Eq. (4) are continuous and can be sampled at arbitrary rates as needed. Particularly, unlike the discrete Fourier transform or discrete wavelet transform, the subspace dimension for 1D occupancy signals $\{f(z)\}$ is not coupled with the sampling grid of $z$-axis. This avoids potential sampling mismatches between the sampling grid used in the training data and that in the testing stage. The reconstruction of $\hat{F}$ from the FOF $\bm{C}$ is sampling scalable in the sense that the sampling rate along the $z$-axis can be chosen as necessary, e.g., according to the requirement of reconstruction quality and speed. 3D models at any resolution can be readily obtained from the FOF representation by simply altering the sampling rate along the $z$-axis.

### v4

@@ I Introduction
Reconstructing a 3D human from a single image has emerged as a popular task in computer vision and graphics, which can be widely used in various downstream applications, such as mixed reality and virtual try-on. However, real-time, high-fidelity monocular 3D human reconstruction remains challenging. The core of this challenge lies in the 3D representation, as it significantly influences the design and performance of reconstruction and generation approaches. Despite some promising results, existing methods [saito2019pifu, sifu, xiu2022icon, econ, liu2025learning, mao2025deep] typically suffer from high computational consumption and lack of robustness. Consequently, a good 3D geometry representation is essential for 3DTV, Holographic Telepresence systems, and other real-time applications, which must fulfill the requirements of accuracy, efficiency, and compatibility [alatan20073DTV, yang2023high].
@@
Classic representations, such as voxel grids [zheng2019deephuman] or meshes [zhu2021HMD, alldieck2019tex2shape, li2021TIP], have been explored for monocular human reconstruction. However, voxel grids require a space complexity of $O(n^{3})$ (where $n$ is the resolution of grids in each dimension), and meshes struggle with topology changes or large deformations. Although some recent methods [econ, tech] propose optimization-based pipelines combining normal maps, voxel grids, and meshes, they all suffer from efficiency issues.
Implicit neural representations have emerged and been widely used in monocular human reconstruction [saito2019pifu, saito2020pifuhd, zheng2021pamir, alldieck2022phorhum, xiu2022icon]. These methods treat 3D space as a continuous field, such as an occupancy or signed distance field, represented as $F(x,y,z):\mathbb{R}^{3}\to\mathbb{R}$. Instead of voxel grids, a neural network models the field, allowing results at any resolution. However, these methods infer values across numerous spatial grid points, which slows down inference and makes high-frame-rate or real-time reconstruction a significant challenge. Based on PIFu [saito2019pifu], Monoport [li2020monoport] proposes an efficient sampling scheme to speed up inference, achieving 15 FPS only for mesh-free rendering. However, when generating a complete mesh, their method operates at less than 10 FPS, and the results are often worse than PIFu. These limitations hinder Monoport’s applicability with downstream applications.
@@
Parametric Model [SMPL:2015, SMPL-X:2019, STAR:2020]
Voxel Grid[zheng2019deephuman]
Mesh[alldieck2019tex2shape, li2021TIP, zhu2021HMD]
Implicit Function[saito2019pifu], [saito2020pifuhd, zheng2021pamir, alldieck2022phorhum, xiu2022icon, sith]
@@
  • Compared with state-of-the-art methods based on other representations, our approach can produce high-quality results in real-time. Our system is the first 30+FPS pipeline, achieving a twofold improvement in speed over Monoport [li2020monoport], coupled with superior quality.
An early version of our method, including only FOF, was published as a conference paper [fof]. Based on our FOF representation, several works for generation[joint2human] and reconstruction[li2024diffusion, yang2024r2human, yang2024real] have been proposed. In this paper, we substantially extend the original version by developing FOF-X. Firstly, we design a new reconstruction framework that greatly mitigates the performance degradation caused by texture and lighting effects. Secondly, we propose a robust mesh-to-FOF conversion algorithm with an automaton-based discontinuity matcher, enabling real-time execution and significantly improving the system’s robustness when facing challenging poses. Thirdly, we propose a FOF-to-mesh algorithm with a Laplacian coordinate constraint for greater robustness and fidelity. Such a strategy effectively addresses artifacts caused by view-direction bias without any loss of geometric details. Fourth, we employ cosine series as the subspace approximation basis, retaining the Fourier-like approximation capacity while eliminating Gibbs phenomena through even-periodic boundary extension. Additionally, we evaluate FOF-X through comprehensive experiments including: testing on more datasets, comparisons with various state-of-the-art methods, and more detailed ablation studies, all confirming the method’s effectiveness, efficiency and robustness.
@@ II Related Work
Monocular 3D human reconstruction methods can be grouped into surface-based, volume-based, and “sandwich-like” approaches. In addition, recent works explore generative models for completion from a single image.
@@ II-A Surface-based Reconstruction
Surface-based approaches focus on inferring the interface between geometry and empty space, and can be broadly classified into three categories: parametric models, UV-map based methods, and graph-based methods. Parametric models such as SMPL [SMPL:2015], SMPL-X [SMPL-X:2019], and STAR [STAR:2020] represent naked human bodies via statistical shape spaces and are widely used to regress meshes from RGB or RGBD images [pymaf2021, hu20213dbodynet]. Some methods [tang2023mlp, yu2023mobirfpose] focus solely on 3D human pose estimation, and their outputs can also be converted into parametric models. However, they cannot capture clothing or hair. UV-map-based methods deform parametric models by estimating per-vertex displacements. Zhu et al[zhu2019HMD1, zhu2021HMD] employ a four-stage process constrained by joints, silhouettes, and shading, but their results are sometimes inconsistent with the images. Graph-based representations are naturally compatible with triangle meshes. Li et al[li2021TIP] employ graph neural networks for topology-consistent reconstructions, but results are smooth and lack details. Habermann et al[Deecap2] capture dense performance from monocular video but require pre-scanned templates. Overall, surface-based methods struggle with topology variation and fine detail recovery.
@@ II-B Volume-based Reconstruction
Volume-based methods predict occupancy or signed distance values, naturally handling arbitrary topology. Early voxel-based CNNs [zheng2019deephuman] are memory-heavy and low-resolution. Implicit representations overcome this, as in PIFu [saito2019pifu] and PIFuHD [saito2020pifuhd], which regress pixel-aligned functions and normals for high-detail geometry. PAMIR [zheng2021pamir] and ICON [xiu2022icon] leverage SMPL priors for robustness, though outputs often misalign with image evidence and appear noisy. Recent works enhance implicit fields with transformers [gta, sifu], but inference is computationally expensive. Li et al[li2020monoport] improve sampling efficiency, yet real-time performance remains elusive. Recent NeRF [2020NeRF] has also been adapted for humans [sherf, elicit], enabling supervision via volume rendering. While SHERF [sherf] generalizes across humans and ELICIT [elicit] leverages CLIP, NeRF-based approaches mainly excel at view synthesis rather than accurate geometry. Meanwhile, GS-SFS [jiang2024gs] reconstructs humans via Gaussian Splatting but relies on dense view inputs and struggles with recovering fine geometric details.
@@
In recent works [modulingH, gin, econ], 3D objects are modeled using front and back depth maps, resembling a sandwich structure that encloses the inner space. However, depth maps only represent part of the geometric surface and cannot fully reconstruct the object. Therefore, these methods require volume-based post-processing, such as IF-Nets [if-nets] and Poisson surface reconstruction[2006Poisson], to complete the result further. Any-Shot GIN [gin] extends the sandwich-like scheme to novel classes of objects by predicting the depth maps of two sides from a single image and performing shape completion with IF-Nets. ECON [econ] follows a similar approach and extends it to 3D humans. It estimates the normal maps of two sides and converts them into depth maps using an optimization method. The parametric human model prior is also employed in the shape completion stage to “inpaint” the missing geometry. However, this optimization process is time-consuming, and its instability may lead to incorrect geometric estimates.
@@ II-D Leveraging Generative Models
Generative models address missing geometry by synthesizing plausible invisible regions. SiTH [sith] generates back views with diffusion to estimate dual-sided normals. TeCH [tech] integrates DMTet with SDS [poole2022dreamfusion] for SMPL-X mesh recovery, yielding visually appealing but noisy surfaces. Human-LRM [humanlrm2023] combines NeRF-based LRM with diffusion-based refinement, though geometric gains are limited. PSHuman [li2025pshuman] produces multi-view images via diffusion and refines SMPL-X meshes with differentiable rasterization, but its heavy pipeline hampers efficiency. Text2Avatar [kwon2025text2avatar] generates avatars from text instructions but still requires dense view inputs for high-quality results. In general, generative methods are time-consuming and often fail to preserve input fidelity.
To solve the problems with existing representations, we propose the Fourier Occupancy Field (FOF), a representation that is both image-aligned and efficient. Building upon FOF, we design FOF-X and parallel inter-conversion algorithms, achieving robust high-quality monocular human reconstruction at 30 FPS.
@@ III-A1 Vanilla FOF
Given a normalized cube $1,1]^{3}$, the occupancy field $F:[-1,1]^{3}\mapsto{0,0.5,1}$ is defined as
@@
Such a 3D representation is highly redundant, because only a small subset, i.e., the iso-surface of $F(x,y,z)$ with value of 0.5, is sufficient to represent the surface of the object. Direct inference of the 3D occupancy field $F$ from a single image not only confronts with the curse of dimensionality, but also requires more computation and memory in handling high dimensional feature maps. Note that the occupancy field defined in the 3D cube can be regarded as the collection of a large number of 1D signals defined on the lines along the view direction. Without loss of generality, we assume that the view direction is the same as the $z$-axis of the 3D cube. The occupancy function, as a 1D signal $f(z):[-1,1]\mapsto\left\{0,0.5,1\right\}$, along the line passing through a particular point $(x^{*},y^{*})\in[-1,1]^{2}$ on the $xy$-plane can be written as $f(z)=F(x^{*},y^{*},z)$. We explore Fourier representation for such a family of 1D occupancy signals and propose a more compact 2D vector field orthogonal to the view direction, namely Fourier Occupancy Field (FOF), for more efficient representation of the 3D occupancy field, which is detailed in the following subsections.
Fourier Series on a Single Occupancy Line: 1D occupancy signals $\left\{f(z)\right\}$ are essentially on-off signals switching at the boundaries of human bodies, which lie in a low-dimensional manifold in the ambient signal space. The $f_{p}(z)$ can be expanded as a convergent Fourier series:
@@
where $\{a_{n}\},\{b_{n}\}\in\mathbb{R}$ are coefficients of basis functions $\left\{\cos(nx)\right\}$ and $\left\{\sin(nx)\right\}$, respectively. Note that by defining the occupancy function as Eq. (1), Eq. (2) holds for discontinuity at $z^{*}$ on the surface $S$ because $f(z^{*})=(f(z^{*}-)+f(z^{*}+))/2=0.5$, where $f(z^{*}-)$ and $f(z^{*}+)$ are the left hand limit and right hand limit of $f(z)$ at $z^{*}$. For compact representation, we approximate 1D occupancy signals $\hat{f}(z)$ by a subspace spanned by the first $2N+1$ basis functions:
@@
Fourier Occupancy Field: Such a Fourier subspace approximation is applied to all 1D occupancy signals over the $xy$-plane:
@@
where $\bm{C}(x,y)$ is the ($2N+1$)-dimensional Fourier coefficient vector for the 1D occupancy signal at $(x,y)$. In this way, we obtain the Fourier Occupancy Field $\bm{C}:[-1,1]^{2}\mapsto\mathbb{R}^{2N+1}$ for 3D occupancy field $\hat{F}(x,y,z)$.